# Nonlinear Schrödinger-Poisson systems in dimension two: the zero mass case

### v1

@@ 4.1. Mountain pass level estimate in the critical case
Proof.
Since $\{u_{n}\}_{n}$ is bounded in the reflexive space $X$, there exists $u\in X$ such that $u_{n}\rightharpoonup u$ in $X$, which implies and that, up to a subsequence, the convergence is strong in $L^{q}(\mathbb{R}^{2})$, for $q\geq p$, and a.e. in $\mathbb{R}^{2}$ thanks to Lemma 2.1. Let us define the function
(4.35) $G(t):=\int_{0}^{t}\sqrt{\frac{F(s)f^{\prime}(s)}{f(s)^{2}}}\,{\rm d}s\,,$
which is well-defined and $C^{1}$ thanks to ($f_{2}$), and let $v_{n}:=G(u_{n})$. Then, by ($f_{2}$) one has
and, by $(f_{1})$ and Hölder’s inequality,
$\begin{split}|G(t)|^{p}&\leq t^{\tfrac{p}{2}}\left(\int_{0}^{t}\frac{F(s)f^{\prime}(s)}{f(s)^{2}}\,{\rm d}s\right)^{\frac{p}{2}}\leq t^{\tfrac{p}{2}}\left(-\frac{F(t)}{f(t)}+\lim_{s\to 0^{+}}\frac{F(s)}{f(s)}+t\right)^{\frac{p}{2}}\leq t^{p}\left(-\frac{F(t)}{tf(t)}+1\right)^{\frac{p}{2}}\leq t^{p},\end{split}$
which yields at once $\|v_{n}\|_{p}\leq\|u_{n}\|_{p}$ and $\|v_{n}\|_{*,p}\leq\|u_{n}\|_{*,p}$. Note also that $G(t)>\sqrt{\tau}t$, hence $v_{n}>\sqrt{\tau}\,u_{n}$ a.e. in $\mathbb{R}^{2}$. Recalling now that $u_{n}\geq 0$, and combining (4.17) with $\mathcal{I}(u_{n})\to c_{mp}$, we infer
$\begin{split}2c_{mp}+\text{o}_{n}(1)&=\|\nabla u_{n}\|_{2}^{2}+\int_{\mathbb{R}^{2}}\left(\ln|\cdot|\ast F(u_{n})\right)F(u_{n})\\ &=\|\nabla u_{n}\|_{2}^{2}-\int_{\mathbb{R}^{2}}|\nabla u_{n}|^{2}\left(1-\frac{F(u_{n})f^{\prime}(u_{n})}{f(u_{n})^{2}}\right)\\ &=\int_{\mathbb{R}^{2}}|\nabla u_{n}|^{2}\left(\frac{F(u_{n})f^{\prime}(u_{n})}{f(u_{n})^{2}}\right)=\|\nabla v_{n}\|_{2}^{2}\,.\end{split}$
Since $c_{mp}<\frac{2\pi}{\alpha_{0}}$ by assumption, we deduce
(4.36) $\|\nabla v_{n}\|_{2}^{2}=2c_{mp}+\text{o}_{n}(1)<\frac{4\pi}{\alpha_{0}}\,.$
Moreover, by ($f_{3}$), as in [14, Lemma 6.3], for all $\varepsilon>0$ there exists $t_{\varepsilon}>0$ such that
$u_{n}\leq t_{\varepsilon}+\frac{v_{n}}{1-\varepsilon}\quad\ \mbox{for all}\ \ x\in\mathbb{R}^{2}.$
From all this, together with (2.7) with $\alpha>\alpha_{0}$, we may estimate as follows:
$\begin{split}\int_{\mathbb{R}^{2}}|F(u_{n})|^{\gamma}&\leq C_{\varepsilon}\int_{\{u_{n}<t_{\varepsilon}\}}|u_{n}|^{p\gamma}+C_{\varepsilon}\int_{\{u_{n}\geq t_{\varepsilon}\}}\left(t_{\varepsilon}+\frac{v_{n}}{1-\varepsilon}\right)^{\gamma r}{\rm e}^{\gamma\alpha(1+\varepsilon)\left(t_{\varepsilon}+\frac{v_{n}}{1-\varepsilon}\right)^{2}}\,{\rm d}x\\ &\leq C_{\varepsilon}\int_{\mathbb{R}^{2}}|u_{n}|^{p\gamma}+C_{\varepsilon}\int_{\{u_{n}\geq t_{\varepsilon}\}}{\rm e}^{\gamma\alpha\left(\frac{1+\varepsilon}{1-\varepsilon}\right)^{2}\|\nabla v_{n}\|_{2}^{2}\left(\frac{v_{n}}{\|\nabla v_{n}\|_{2}}\right)^{2}}\,{\rm d}x\,,\end{split}$
where in the last step we used the inequalities $s^{r}\leq C_{\varepsilon}\,{\rm e}^{\varepsilon s^{2}}$ for any $s\geq t_{\varepsilon}$ and
$\left(t_{\varepsilon}+\frac{v_{n}}{1-\varepsilon}\right)^{2}\leq C_{\varepsilon}t_{\varepsilon}^{2}+(1+\varepsilon)\left(\frac{v_{n}}{1-\varepsilon}\right)^{2},$
see [14, eq. (6.9)]. Choosing now $\varepsilon$ small enough such that $t_{\varepsilon}\geq\tau^{-1/2}$, in the set $\{u_{n}\geq t_{\varepsilon}\}$ one has $v_{n}\geq 1$, and therefore, defining $w_{n}:=\tfrac{v_{n}}{\|\nabla v_{n}\|_{2}}$, by means of the (strict) fine upper bound in (4.36), Theorem 2.2 with $q=p$, $\gamma\in[1,\frac{2\pi}{\alpha_{0}c_{mp}})$, and choosing $\alpha$ close to $\alpha_{0}$, one has
(4.37) $\begin{split}\int_{\mathbb{R}^{2}}|F(u_{n})|^{\gamma}&\leq C_{\varepsilon}\int_{\mathbb{R}^{2}}|u_{n}|^{p\gamma}+C_{\varepsilon}\int_{\mathbb{R}^{2}}|v_{n}|^{p}{\rm e}^{\gamma\alpha\left(\frac{1+\varepsilon}{1-\varepsilon}\right)^{2}\|\nabla v_{n}\|_{2}^{2}\left(\frac{v_{n}}{\|\nabla v_{n}\|_{2}}\right)^{2}}\,{\rm d}x\\ &\leq C_{\varepsilon}\int_{\mathbb{R}^{2}}|u_{n}|^{p\gamma}+C_{\varepsilon}\|\nabla v_{n}\|_{2}^{p}\int_{\mathbb{R}^{2}}w_{n}^{p}{\rm e}^{\gamma\alpha\left(\frac{1+\varepsilon}{1-\varepsilon}\right)^{2}\|\nabla v_{n}\|_{2}^{2}w_{n}^{2}}\,{\rm d}x\\ &\leq C_{\varepsilon}\|u_{n}\|_{p\gamma}^{p\gamma}+C_{\varepsilon}\|v_{n}\|_{p}^{p}\\ &\leq C_{\varepsilon}\|u_{n}\|_{p\gamma}^{p\gamma}+C_{\varepsilon}\|u_{n}\|_{p}^{p}\leq C\,,\end{split}$
since $\{u_{n}\}_{n}$ is bounded in $X$ and by Lemma 2.1. ∎
From the proof of Lemma 4.6, and in particular combining (4.37) with the compact embedding provided by Lemma 2.1, we can say actually more:
Corollary 4.7.
Assume $(f_{1})-(f_{4})$ and $(f_{c})$ hold. Let $\{u_{n}\}_{n}\subset X$ be a nonnegative Cerami sequence which is bounded in $X$, and such that $u_{n}\to 0$ in $L^{s}(\mathbb{R}^{2})$ for all $s>p$. Then, for any $\gamma\in\left(1,\frac{2\pi}{\alpha_{0}c_{mp}}\right)$
@@
Reasoning in the same way as for (4.37) in Lemma 4.6, for some $r>0$ and $\alpha>\alpha_{0}$ we get
$\int_{\mathbb{R}^{2}}|F(u_{n})|^{\gamma}\,{\rm d}x\leq C_{\varepsilon}\int_{\mathbb{R}^{2}}|u_{n}|^{p\gamma}+C_{\varepsilon}\int_{\mathbb{R}^{2}}|v_{n}|^{r}{\rm e}^{\gamma\alpha\left(\frac{1+\varepsilon}{1-\varepsilon}\right)^{2}\|\nabla v_{n}\|_{2}^{2}\left(\frac{v_{n}}{\|\nabla v_{n}\|_{2}}\right)^{2}}\!\,{\rm d}x\,.$
The first term on the right-hand side goes to 00 as $n\to+\infty$ since $p\gamma>p$, while for the second term we proceed by using Hölder’s inequality and Theorem 2.2: indeed, choosing $\sigma\in(1,\sigma_{0})$, with $\sigma_{0}$ close to $1$ such that the exponent is less than $4\pi$, $r$ big enough such that $\frac{r}{2}\sigma>p$, and $\alpha$ close to $\alpha_{0}$, we obtain
$\int_{\mathbb{R}^{2}}|F(u_{n})|^{\gamma}\,{\rm d}x$ $\leq C_{\varepsilon}\|u_{n}\|^{p\gamma}_{p\gamma}+C_{\varepsilon}\left(\int_{\mathbb{R}^{2}}|v_{n}|^{\frac{r}{2}\sigma^{\prime}}\,{\rm d}x\right)^{\frac{1}{\sigma^{\prime}}}\left(\int_{\mathbb{R}^{2}}|v_{n}|^{\frac{r}{2}\sigma}{\rm e}^{\sigma\gamma\alpha\left(\frac{1+\varepsilon}{1-\varepsilon}\right)^{2}\|\nabla v_{n}\|_{2}^{2}\left(\frac{v_{n}}{\|\nabla v_{n}\|_{2}}\right)^{2}}\,{\rm d}x\right)^{\frac{1}{\sigma}}$
$\leq C_{\varepsilon}\|u_{n}\|^{p\gamma}_{p\gamma}+C_{\varepsilon}\|u_{n}\|^{\frac{r}{2}}_{\frac{r}{2}\sigma^{\prime}}C\|v_{n}\|^{\frac{r}{2}\sigma}_{\frac{r}{2}\sigma}\,,$
which tends to 00 as $n\to+\infty$ since $p\gamma>s$.
Concerning $\|f(u_{n})u_{n}\|_{\gamma}$, we can repeat the same computations as above, using now $(f_{1})$ and $(f_{c})$, hence
$\int_{\mathbb{R}^{2}}|f(u_{n})u_{n}|^{\gamma}\,{\rm d}x\to 0$
as $n\to+\infty$, concluding the proof. ∎
@@
By Lemma 4.1 we know that $\mathcal{I}$ satisfies the mountain pass geometry, and this yields the existence of a Cerami sequence $\{u_{n}\}_{n}\subset X$ for $\mathcal{I}$ at level $c_{mp}$ defined in (4.3), satisfying (4.14) by Lemma 4.2, which can be assumed nonnegative by Remark 4.4. By Lemma 4.3 such a sequence is bounded in $X$. Suppose by contradiction that $\{u_{n}\}_{n}$ is vanishing, namely
@@
Since $\{u_{n}\}_{n}$ is bounded, then $u_{n}\to 0$ in $L^{s}(\mathbb{R}^{2})$ as $n\to+\infty$ for every $s\in(p,+\infty)$ by [32, Lemma I.1]. By exploiting the inequality $\ln\left(1+\tfrac{b}{t}\right)\lesssim t^{q}$ for $q\in(0,1]$, and the Hardy-Littlewood-Sobolev inequality, if ($f_{c}$) is assumed one may estimate
(5.2) $\begin{split}I_{2}(u_{n})&=\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}\ln\left(1+\frac{b}{|x-y|}\right)F(u_{n}(x))F(u_{n}(y))\,{\rm d}x\,{\rm d}y\\ &\lesssim\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}\frac{F(u_{n}(x))F(u_{n}(y))}{|x-y|^{\frac{4(\gamma-1)}{\gamma}}}\,{\rm d}x\,{\rm d}y\lesssim\|F(u_{n})\|_{\gamma}^{2}\to 0\end{split}$
as $n\to+\infty$, which holds by Corollary 4.7 for $\gamma\in\left(1,\min\left\{\frac{4}{3},\frac{2\pi}{\alpha_{0}c_{mp}}\right\}\right]$. On the other hand, when ($f_{sc}$) holds, by combining (2.6) with $r=p/2$ and Theorem 2.2 we get
$\begin{split}\int_{\mathbb{R}^{2}}|F(u_{n})|^{\gamma}&\lesssim\int_{\mathbb{R}^{2}}|u_{n}|^{\gamma p}+\int_{\mathbb{R}^{2}}|u_{n}|^{\gamma\frac{p}{2}}\,{\rm e}^{\alpha|u_{n}|^{2}}\\ &\leq\|u_{n}\|_{\gamma p}+\|u_{n}\|_{\gamma p}^{\gamma\frac{p}{2}}\left(\int_{\mathbb{R}^{2}}|u_{n}|^{\gamma p}\,{\rm e}^{2\alpha u_{n}^{2}}\right)^{\frac{1}{2}}\\ &\leq\|u_{n}\|_{\gamma p}+\|u_{n}\|_{\gamma p}^{\gamma\frac{p}{2}}\|\nabla u_{n}\|_{2}^{(\gamma-1)p}\|u_{n}\|_{p}^{p}\to 0\,,\end{split}$
@@
(5.3) $A_{2}(F(u_{n}),f(u_{n})u_{n})=\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}\ln\left(1+\frac{b}{|x-y|}\right)F(u_{n}(x))f(u_{n}(y))u_{n}(y)\,{\rm d}x\,{\rm d}y\to 0$
@@
(5.4) $\begin{split}2c_{mp}+\text{o}_{n}(1)&=2\mathcal{I}(u_{n})-\mathcal{I}^{\prime}(u_{n})[u_{n}]=I_{0}(u_{n})-A_{0}(F(u_{n}),f(u_{n})u_{n})\\ &=2I_{1}(u_{n})-A_{1}(F(u_{n}),f(u_{n})u_{n})-2I_{2}(u_{n})+A_{2}(F(u_{n}),f(u_{n})u_{n})\\ &=\int_{\mathbb{R}^{2}}\left(\ln(b+|\cdot|)\ast F(u_{n})\right)\left(F(u_{n})-f(u_{n})u_{n}\right)\,{\rm d}x+\text{o}_{n}(1)<0\end{split}$
@@
by (2.8). Since $\{\widetilde{u}_{n}\}_{n}$ is hence bounded in $X$, there exists $\widetilde{u}\in X$ such that $\widetilde{u}_{n}\rightharpoonup\widetilde{u}$ in $X$ and $\widetilde{u}_{n}\to\widetilde{u}$ in $L^{q}(\mathbb{R}^{2})$ for all $q\geq p$, as well as a.e. in $\mathbb{R}^{2}$ by Lemma 2.1. By (5.5) it is easy to see that
(5.8) $\delta<\|\widetilde{u}_{n}\|_{L^{p}(B_{R}(0))}^{p}\to\|\widetilde{u}\|_{L^{p}(B_{R}(0))}^{p}\,,$
@@
(5.9) $\|\varphi(\cdot-y_{n})\|^{2}=\|\nabla\varphi\|_{2}^{2}+\int_{\mathbb{R}^{2}}\ln(b+|x+y_{n}|)|\varphi(x)|^{p}\,{\rm d}x\leq\|\varphi\|^{2}+\ln(b+|y_{n}|)\|\varphi\|_{p}^{p}\,.$
@@
(5.10) $\begin{split}\|u_{n}\|_{\ast,p}^{p}&=\int_{\mathbb{R}^{2}}\ln(b+|x+y_{n}|)|\widetilde{u}_{n}(x)|^{p}\,{\rm d}x\\ &\geq\int_{B_{1}(0)}\ln(b+|x+y_{n}|)|\widetilde{u}_{n}(x)|^{p}\,{\rm d}x\geq\ln|y_{n}|\int_{B_{1}(0)}|\widetilde{u}_{n}|^{p}\\ &\geq\delta\ln|y_{n}|\geq\delta\ln(b+|y_{n}|)\frac{\ln b}{\ln(2b)}\,.\end{split}$
@@
$\begin{split}|\mathcal{I}^{\prime}(\widetilde{u}_{n})[\varphi]|&\leq\|\mathcal{I}^{\prime}(u_{n})\|_{X^{\prime}}\left[\|\varphi\|^{2}+\ln(2b)\left(\frac{1}{\delta\ln b}\|u_{n}\|_{\ast,p}^{p}+1\right)\|\varphi\|_{p}^{p}\right]^{\frac{1}{2}}\\ &\leq C\|\mathcal{I}^{\prime}(u_{n}))\|_{X^{\prime}}\left[\|\varphi\|^{2}+\|\varphi\|_{p}^{p}\right]\to 0\,,\end{split}$
as $n\to+\infty$, by recalling that $\|u_{n}\|_{\ast,p}$ is uniformly bounded and that $\|\mathcal{I}^{\prime}(u_{n})\|_{X^{\prime}}\to 0$ by Lemma 4.3.
Since $\mathcal{I}^{\prime}(\widetilde{u}_{n})\to 0$ in $X^{\prime}$ and $\mathcal{I}^{\prime}(\widetilde{u}_{n})[\widetilde{u}_{n}]\to 0$, we have then
(5.11) $\text{o}_{n}(1)=\mathcal{I}^{\prime}(\widetilde{u}_{n})[\widetilde{u}_{n}-\widetilde{u}]=\int_{\mathbb{R}^{2}}\nabla\widetilde{u}_{n}\nabla(\widetilde{u}_{n}-\widetilde{u})+A_{0}(F(\widetilde{u}_{n}),f(\widetilde{u}_{n})(\widetilde{u}_{n}-\widetilde{u}))\,.$
@@
(5.12) $\int_{\mathbb{R}^{2}}\nabla\widetilde{u}_{n}\nabla(\widetilde{u}_{n}-\widetilde{u})=\int_{\mathbb{R}^{2}}|\nabla(\widetilde{u}_{n}-\widetilde{u})|^{2}+\int_{\mathbb{R}^{2}}\nabla\widetilde{u}\,\nabla(\widetilde{u}_{n}-\widetilde{u})=\|\widetilde{u}_{n}-\widetilde{u}\|^{2}_{D^{1,2}(\mathbb{R}^{2})}+\text{o}_{n}(1)$
since $\widetilde{u}_{n}\rightharpoonup\widetilde{u}$ in $D^{1,2}(\mathbb{R}^{2})$. Moreover, as in (5.2), one infers

### v2

@@ 4.1. Mountain pass level estimate in the critical case
If, moreover, $u_{n}\to 0$ in $L^{s}(\mathbb{R}^{2})$ for all $s>p$, then for any $\gamma\in\left(1,\frac{2\pi}{\alpha_{0}c_{mp}}\right)$ one has
@@
Since $\{u_{n}\}_{n}$ is bounded in the reflexive space $X$, there exists $u\in X$ such that $u_{n}\rightharpoonup u$ in $X$, which implies that, up to a subsequence, the convergence is strong in $L^{q}(\mathbb{R}^{2})$, for $q\geq p$, and a.e. in $\mathbb{R}^{2}$ thanks to Proposition 2.1. From Lemma 4.6, together with (2.7) with $\alpha>\alpha_{0}$, we may estimate as follows:
$\int_{\mathbb{R}^{2}}|F(u_{n})|^{\gamma}$ $\leq C_{\varepsilon}\int_{\{u_{n}<t_{\varepsilon}\}}|u_{n}|^{p\gamma}+C_{\varepsilon}\int_{\{u_{n}\geq t_{\varepsilon}\}}\left(t_{\varepsilon}+\frac{v_{n}}{1-\varepsilon}\right)^{\gamma r}{\rm e}^{\gamma\alpha(1+\varepsilon)\left(t_{\varepsilon}+\frac{v_{n}}{1-\varepsilon}\right)^{2}}$
$\leq C_{\varepsilon}\int_{\mathbb{R}^{2}}|u_{n}|^{p\gamma}+C_{\varepsilon}\int_{\{u_{n}\geq t_{\varepsilon}\}}{\rm e}^{\gamma\alpha\left(\frac{1+\varepsilon}{1-\varepsilon}\right)^{2}\|\nabla v_{n}\|_{2}^{2}\left(\frac{v_{n}}{\|\nabla v_{n}\|_{2}}\right)^{2}}\,,$
where in the last step we used $s^{\gamma r}\leq C_{\varepsilon}\,{\rm e}^{\varepsilon s^{2}}$ for any $s\geq t_{\varepsilon}$ and
(4.35) $\left(t_{\varepsilon}+\frac{v_{n}}{1-\varepsilon}\right)^{2}\leq C_{\varepsilon}t_{\varepsilon}^{2}+(1+\varepsilon)\left(\frac{v_{n}}{1-\varepsilon}\right)^{2},$
by the $\varepsilon-$Young inequality. Choosing now $\varepsilon$ small enough such that $t_{\varepsilon}\geq\tau^{-1/2}$, in the set $\{u_{n}\geq t_{\varepsilon}\}$ one has $v_{n}\geq 1$, and therefore, by means of the (strict) fine upper bound in (4.33), Theorem 2.2 with $q=p$, $\gamma\in[1,\frac{2\pi}{\alpha_{0}c_{mp}})$, and choosing $\alpha$ close to $\alpha_{0}$ and $\varepsilon$ small, one gets
(4.36) $\int_{\mathbb{R}^{2}}|F(u_{n})|^{\gamma}$ $\leq C_{\varepsilon}\int_{\mathbb{R}^{2}}|u_{n}|^{p\gamma}+C_{\varepsilon}\int_{\mathbb{R}^{2}}|v_{n}|^{p}{\rm e}^{\gamma\alpha\left(\frac{1+\varepsilon}{1-\varepsilon}\right)^{2}\|\nabla v_{n}\|_{2}^{2}\left(\frac{v_{n}}{\|\nabla v_{n}\|_{2}}\right)^{2}}$
$\leq C_{\varepsilon}\int_{\mathbb{R}^{2}}|u_{n}|^{p\gamma}+C_{\varepsilon}\|\nabla v_{n}\|_{2}^{p}\int_{\mathbb{R}^{2}}\left|\frac{v_{n}}{\|\nabla v_{n}\|_{2}}\right|^{p}{\rm e}^{\gamma\alpha\left(\frac{1+\varepsilon}{1-\varepsilon}\right)^{2}\|\nabla v_{n}\|_{2}^{2}\left(\tfrac{v_{n}}{\|\nabla v_{n}\|_{2}}\right)^{2}}$
$\leq C_{\varepsilon}\|u_{n}\|_{p\gamma}^{p\gamma}+C_{\varepsilon}\|v_{n}\|_{p}^{p}\leq C_{\varepsilon}\|u_{n}\|_{p\gamma}^{p\gamma}+C_{\varepsilon}\|u_{n}\|_{p}^{p}\leq C\,,$
since $\{u_{n}\}_{n}$ is bounded in $X$ and by Proposition 2.1.
Suppose now in addition that $u_{n}\to 0$ in $L^{s}(\mathbb{R}^{2})$ for all $s>p$. In the same way as in (4.36) for some $r>0$ and $\alpha>\alpha_{0}$ one has
$\int_{\mathbb{R}^{2}}|F(u_{n})|^{\gamma}$ $\leq C_{\varepsilon}\int_{\mathbb{R}^{2}}|u_{n}|^{p\gamma}+C_{\varepsilon}\int_{\mathbb{R}^{2}}|v_{n}|^{r\gamma}{\rm e}^{\gamma\alpha\left(\frac{1+\varepsilon}{1-\varepsilon}\right)^{2}\|\nabla v_{n}\|_{2}^{2}\left(\frac{v_{n}}{\|\nabla v_{n}\|_{2}}\right)^{2}}\!\,{\rm d}x$
$\quad\leq C_{\varepsilon}\|u_{n}\|^{p\gamma}_{p\gamma}+C_{\varepsilon}\|v_{n}\|_{\frac{r}{2}\sigma^{\prime}\gamma}^{\frac{r}{2}\gamma}\left(\int_{\mathbb{R}^{2}}|v_{n}|^{\frac{r}{2}\sigma\gamma}{\rm e}^{\sigma\gamma\alpha\left(\frac{1+\varepsilon}{1-\varepsilon}\right)^{2}\|\nabla v_{n}\|_{2}^{2}\left(\frac{v_{n}}{\|\nabla v_{n}\|_{2}}\right)^{2}}\,{\rm d}x\right)^{\frac{1}{\sigma}}.$
The first term on the right-hand side goes to 00 as $n\to+\infty$ since $p\gamma>p$; for the second term we proceed by using Hölder’s inequality and Theorem 2.2: indeed, choosing $\alpha$ close to $\alpha_{0}$, $\sigma>1$ close to $1$, and $\varepsilon>0$ small, since $\gamma<\frac{2\pi}{c_{mp}\alpha_{0}}$ and by (4.33) the exponent is less than $4\pi$; if we then fix $r>2p$, we obtain
$\int_{\mathbb{R}^{2}}|F(u_{n})|^{\gamma}$ $\lesssim\|u_{n}\|^{p\gamma}_{p\gamma}+\|v_{n}\|^{\frac{r}{2}\sigma^{\prime}\gamma}_{\frac{r}{2}\gamma}\|\nabla v_{n}\|_{2}^{\frac{r}{2}\gamma}\lesssim\|u_{n}\|^{p\gamma}_{p\gamma}+\|v_{n}\|^{\frac{r}{2}\sigma^{\prime}\gamma}_{\frac{r}{2}\gamma}\to 0$
as $n\to+\infty$ since $\frac{r}{2}\gamma>p\gamma>p$. In the last step we used again (4.33).
The proof of $\|f(u_{n})u_{n}\|_{\gamma}\to 0$ follows the same line using (2.8). ∎
@@
By Lemma 4.1 we know that $\mathcal{I}$ satisfies the mountain pass geometry. This yields the existence of a Cerami sequence $\{u_{n}\}_{n}\subset X$ for $\mathcal{I}$ at level $c_{mp}$ defined in (4.3), satisfying (4.14) by Lemma 4.2, and which can be assumed nonnegative by Remark 4.4. By Lemma 4.3 such a sequence is bounded in $X$. Suppose by contradiction that $\{u_{n}\}_{n}$ is vanishing, namely
@@
Since $\{u_{n}\}_{n}$ is bounded, then $u_{n}\to 0$ in $L^{s}(\mathbb{R}^{2})$ as $n\to+\infty$ for every $s\in(p,+\infty)$ by [33, Lemma I.1]. By exploiting the inequality $\ln\left(1+\tfrac{b}{t}\right)\lesssim t^{-q}$ for $q\in(0,1]$ and $t\in\mathbb{R}$, and the Hardy-Littlewood-Sobolev inequality, if ($f_{c}$) is assumed one may estimate
(5.2) $I_{2}(u_{n})$ $=\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}\ln\left(1+\frac{b}{|x-y|}\right)F(u_{n}(x))F(u_{n}(y))\,{\rm d}x\,{\rm d}y$
$\lesssim\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}\frac{F(u_{n}(x))F(u_{n}(y))}{|x-y|^{\frac{4(\gamma-1)}{\gamma}}}\,{\rm d}x\,{\rm d}y\lesssim\|F(u_{n})\|_{\gamma}^{2}\to 0$
as $n\to+\infty$, which holds for $\gamma\in\left(1,\min\left\{\frac{4}{3},\frac{2\pi}{\alpha_{0}c_{mp}}\right\}\right)$ by Lemma 4.7. On the other hand, when ($f_{sc}$) holds, by combining (2.7) with $r=p/2$ and Theorem 2.2 with $\alpha$ sufficiently small, we get
(5.3) $\int_{\mathbb{R}^{2}}|F(u_{n})|^{\gamma}$ $\lesssim\int_{\mathbb{R}^{2}}|u_{n}|^{\gamma p}+\int_{\mathbb{R}^{2}}|u_{n}|^{\gamma\frac{p}{2}}\,{\rm e}^{\gamma\alpha|u_{n}|^{2}}$
$\leq\|u_{n}\|_{\gamma p}^{\gamma p}+\|u_{n}\|_{\gamma p}^{\gamma\frac{p}{2}}\left(\int_{\mathbb{R}^{2}}|u_{n}|^{\gamma p}\,{\rm e}^{2\gamma\alpha u_{n}^{2}}\right)^{\frac{1}{2}}$
$\leq\|u_{n}\|_{\gamma p}^{\gamma p}+\|u_{n}\|_{\gamma p}^{\gamma\frac{p}{2}}\|\nabla u_{n}\|_{2}^{(\gamma-1)p}\|u_{n}\|_{p}^{p}\to 0\,,$
@@
$A_{2}(F(u_{n}),f(u_{n})u_{n})=\int_{\mathbb{R}^{2}}\int_{\mathbb{R}^{2}}\ln\left(1+\frac{b}{|x-y|}\right)F(u_{n}(x))f(u_{n}(y))u_{n}(y)\,{\rm d}x\,{\rm d}y\to 0$
@@
(5.4) $2c_{mp}+\text{o}_{n}(1)$ $=2\mathcal{I}(u_{n})-\mathcal{I}^{\prime}(u_{n})[u_{n}]=I_{0}(u_{n})-A_{0}(F(u_{n}),f(u_{n})u_{n})$
$=2I_{1}(u_{n})-A_{1}(F(u_{n}),f(u_{n})u_{n})-2I_{2}(u_{n})+A_{2}(F(u_{n}),f(u_{n})u_{n})$
$=\int_{\mathbb{R}^{2}}\left(\ln(b+|\cdot|)\ast F(u_{n})\right)\left(F(u_{n})-f(u_{n})u_{n}\right)\,{\rm d}x+\text{o}_{n}(1)<0$
@@
by (2.9). Since $\{\widetilde{u}_{n}\}_{n}$ is hence bounded in $X$, there exists $\widetilde{u}\in X$ such that $\widetilde{u}_{n}\rightharpoonup\widetilde{u}$ in $X$ and $\widetilde{u}_{n}\to\widetilde{u}$ in $L^{q}(\mathbb{R}^{2})$ for all $q\geq p$ as well as a.e. in $\mathbb{R}^{2}$ by Proposition 2.1. By (5.5) it is easy to see that
(5.8) $\delta<\|\widetilde{u}_{n}\|_{L^{p}(B_{1})}^{p}\to\|\widetilde{u}\|_{L^{p}(B_{1})}^{p}\,,$
@@
(5.9) $\|\varphi(\cdot-y_{n})\|^{2}\leq\|\varphi\|^{2}+\ln(b+|y_{n}|)\|\varphi\|_{p}^{p}\,.$
@@
(5.10) $\|u_{n}\|_{\ast,p}^{p}$ $\geq\int_{B_{1}}\ln(b+|x+y_{n}|)|\widetilde{u}_{n}(x)|^{p}\,{\rm d}x\geq\ln|y_{n}|\int_{B_{1}}|\widetilde{u}_{n}|^{p}$
$\geq\delta\ln|y_{n}|\geq\delta\ln(b+|y_{n}|)\frac{\ln b}{\ln(2b)}\,.$
@@
$|\mathcal{I}^{\prime}(\widetilde{u}_{n})[\varphi]|$ $\leq\|\mathcal{I}^{\prime}(u_{n})\|_{X^{\prime}}\left[\|\varphi\|^{2}+\ln(2b)\left(\frac{1}{\delta\ln b}\|u_{n}\|_{\ast,p}^{p}+1\right)\|\varphi\|_{p}^{p}\right]^{\frac{1}{2}}$
$\leq C\|\mathcal{I}^{\prime}(u_{n})\|_{X^{\prime}}\left[\|\varphi\|^{2}+\|\varphi\|_{p}^{p}\right]\to 0\,,$
as $n\to+\infty$, by recalling that $\|\mathcal{I}^{\prime}(u_{n})\|_{X^{\prime}}\to 0$ and that $\|u_{n}\|_{\ast,p}$ is uniformly bounded by Lemma 4.3.
Since then $\mathcal{I}^{\prime}(\widetilde{u}_{n})\to 0$ in $X^{\prime}$ and $\mathcal{I}^{\prime}(\widetilde{u}_{n})[\widetilde{u}_{n}]\to 0$, we deduce that
(5.11) $\text{o}_{n}(1)=\mathcal{I}^{\prime}(\widetilde{u}_{n})[\widetilde{u}_{n}-\widetilde{u}]=\int_{\mathbb{R}^{2}}\nabla\widetilde{u}_{n}\cdot\nabla(\widetilde{u}_{n}-\widetilde{u})+A_{0}(F(\widetilde{u}_{n}),f(\widetilde{u}_{n})(\widetilde{u}_{n}-\widetilde{u}))\,.$
@@
(5.12) $\int_{\mathbb{R}^{2}}\nabla\widetilde{u}_{n}\cdot\nabla(\widetilde{u}_{n}-\widetilde{u})$ $=\int_{\mathbb{R}^{2}}|\nabla(\widetilde{u}_{n}-\widetilde{u})|^{2}+\int_{\mathbb{R}^{2}}\nabla\widetilde{u}\cdot\nabla(\widetilde{u}_{n}-\widetilde{u})=\|\widetilde{u}_{n}-\widetilde{u}\|^{2}_{D^{1,2}(\mathbb{R}^{2})}+\text{o}_{n}(1)$
since $\widetilde{u}_{n}\rightharpoonup\widetilde{u}$ in $D^{1,2}(\mathbb{R}^{2})$. Moreover, we claim that